# Under the hood of a carbon footprint calculator

### v1

@@ Abstract.
We explain the mathematical theory of the Input-Output method for carbon footprints computations.
Introduction
@@
Two main methods for calculating carbon footprints are the Input-Output (IO) method and the Lifecycle Analysis (LCA) method, also known as Craddle-to-Grave. Both aim to quantify the greenhouse gas (GHG) emissions of a product or service, but they differ in scope, data needs, and level of detail.
@@
The Input-Output method is an economic approach developed in [L], that uses national or regional data on industries and their interdependencies. It calculates the carbon footprint by tracing the flow of goods and services across different sectors, estimating emissions based on economic input-output tables. It relies on aggregated data about economic sectors (e.g., agriculture, manufacturing) and their emissions, and estimates emissions associated with production, supply chains, and services by analyzing the economic connections between industries. This is the commonly accepted method for large-scale, economy-wide assessments is not data-intensive, relies on publicly available economic data. It will lack detail on specific products or processes and miss the emissions from smaller, niche activities.
The Lifecycle Analysis evaluates a product’s environmental impact over its entire lifecycle — from raw material extraction to disposal. It calculates emissions at each stage, considering both direct and indirect impacts. This provides a detailed, product-specific carbon footprint, identifies emissions hotspots and opportunities for reduction. It is however data-intensive and time-consuming, and the total complexity can lead to global inconsistencies, as the under or over counting of secondary emissions are only estimated.
The main difficulty in the concrete application of the Input-Output method is access to the relevant data: a recent Input-Output table (or transaction matrix) and a measure of environmental impact for each sector appearing in the table. The tables are quite available, see for Switzerland [IOCH]. The latter is harder to obtain, especially as it needs to be computed/collected in an homogeneous manner over the different sectors in order to make sense.
This note aims to explain the mathematical theory of the Input-Output method, and is based on the article by J. Kitzes [JK], the talk by J. Steinberger at the Climathics 2022 conference [Clim], and the third author’s talk for the Comission Romande de mathématiques in 2023.
1. A concrete question
@@
To estimate the carbon footprint of a T-shirt for instance, one needs the footprint of the cotton all the way from production of the raw material up to the stitching of the T-shirt, all the washing during the lifespan of the T-shirt, and the disposal of the unusable T-shirt. But to estimate the cotton footprint only, one needs to estimate the transport of the water, the one of the pesticides used to protect the crop, the fertilizing of the field, all those requiring tractors and fuel, the fuel itself requiring transport and fuel, thus creating loops that are difficult to extricate. Moreover, once one estimates all the emissions, there’s no guarantee that all the emissions of all the objects that we so compute, will actually sum up to all the emissions that we actually produce. The Input-Output method takes the problem from the other end: it takes all the measured emissions, and then attributes it to different sectors, and then computes the carbon footprint of a goods from the textile industry according to its direct emissions, combined with all the indirect emissions as we shall now explain.
2. Basic Input-Output
@@
$\begin{array}[]{c|cccc|cc}&S_{1}&S_{2}&\dots&S_{n}&D&T\\ \hline\cr S_{1}&c_{11}&c_{12}&\dots&c_{1n}&d_{1}&t_{1}\\ S_{2}&c_{21}&c_{22}&\dots&c_{2n}&d_{2}&t_{2}\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots\\ S_{n}&c_{n1}&c_{n2}&\dots&c_{nn}&d_{n}&t_{n}\\ \hline\cr V&v_{1}&v_{2}&\dots&v_{n}&\\ T&t_{1}&t_{2}&\dots&t_{n}&\end{array}$
@@
for $i=1,\dots,n$. The value $t_{i}$ represents the total output of the sector $S_{i}$, which is also the total input because the system is a closed circuit. Out of these vectors we define the technical coefficient matrix $A=(a_{ij})$ where
$a_{ij}=\frac{c_{ij}}{t_{j}}.$
For a vector $W=(w_{1},\dots,w_{n})$ we denote by $D_{W}$ the diagonal matrix with entries $w_{1},\dots,w_{n}$ in the diagonal, so that $A=CD_{T}^{-1}$. The scalar product of two vectors $X=(xi,\dots,x_{n})$ and $Y^{t}=(y_{1},\dots,y_{n})$ is given by
$X,Y\right>=\sum_{i=1}^{n}x_{i}y_{i}=XY$
where $X$ is a line vectors, namely $1\times n$ matrix, and $Y$ is a column vector, namely an $n\times 1$ matrix and $XY$ is the usual matrix multiplication so a real number.
3. Footprints in the Input-Output model
Economical activities have all sorts of footprints that are correlated to the income and the cost of things. To compute the footprint we need to have a emission vector $E=(e_{1},\dots,e_{n})$ where $e_{i}$ is the direct emission of sector $S_{i}$. It could be millions of tons of $CO_{2}$, measured or estimated directly above the industries of the sector, or square feet measured by the occupancy of the industries, or tons of plastic trash estimated by the collectivities, etc. Then $|E|=e_{1}+\cdots+e_{n}$ is the total footprint of the economical system. We compute the direct intensity of the print by $F=(f_{1},\dots,f_{n})$ with
@@ 3. Footprints in the Input-Output model
$D,F\right>=\sum_{i=1}^{n}d_{i}f_{i}$
it is the footprint directly attributed to the consumers, but it doesn’t take into account all the intermediate exchanges that the sectors had to create the product. If you buy goods for $k$ dollars to industry $S_{i}$, then $kf_{i}$ indicates the direct footprint of the purchase, but not the total footprint of the goods. That one has to take into account the exchanges the sector $S_{i}$ had from the other ones. To do this, we need to compute the total intensity vector $X=(x_{1},\cdots,x_{n})$ that gives the total footprint of the sector $S_{i}$ by money amount of final consumer demand, given by
@@
where $I$ denotes the $n\times n$ identity matrix (most matrices are invertible so let’s assume that one is).
@@
The footprint is completely attributed to the consumers. Namely, with the above notations, $X,D\right>=|E|$.
@@
[…]
$X,D\right>=\left<F(I-A)^{-1},(I-A)T\right>=FT=\sum_{i=1}^{n}\frac{e_{i}}{t_{i}}t_{i}=|E|$
Remark 1.
Similarly one can define $Y=F(I-A^{t})^{-1}$ such that $Y,V\right>=|E|$ and that vector attributes the print to the added value of the different sector.
4. Limitations
@@
The computation doesn’t take into account the added value $V$, which is what the companies make, and the model attributes the footprint completely to the customers, through the vector $D$. Attributing the footprint to the added value created by the sectors would seem more natural to understand their responsibility, see Remark 1 above.
Inverses
Mathematically speaking, the main problem is that if $B$ is a matrix $\epsilon$-close to $A$ (say in terms of the coefficients), then the distance between $(I-A)^{-1}$ and $(I-B)^{-1}$ could be a priori be as big as we want, for instance if the matrices both have an eigenvalue close enough to 1. Since all the values are estimated, the model is probably not very stable in an optimization exercise.
Life Cycle Analysis (LCA)
To get a comprehensive and detailed view of the total carbon footprint of a given goods, one can examine each stage, including production, transportation, usage, and end-of-life processes. This method can estimate the footprint of goods with a detail that Input-Output can’t, but uses data-intensive estimates that need to be computed, with not total consistency guarantee whatsoever, and no way to estimate the inconsistency.

### v2

@@ Abstract.
We explain the basic linear algebra formulas used in carbon footprints computations.
@@
Have you computed your carbon footprint lately? Wondered where all those numbers came from? Two main methods for calculating carbon footprints are the Input-Output method and the Lifecycle Analysis method, also known as Craddle-to-Grave. Both aim to quantify the emissions of a product or service, but the methods are fundamentally different.
@@
Input-Output is an economic approach developed in [Leo], widely accepted since the 60ies, but only recently adapted to environmental reporting. It uses national or regional data on industries and their inter-dependencies. It relies on aggregated data about economic sectors (e.g., agriculture, manufacturing) and their emissions, and we shall see in this note (Definition 2) the formula estimating emissions associated with production, supply chains, and services. This is the commonly accepted method for large-scale, economy-wide assessments, is not data-intensive, relies on publicly available economic data. It will lack detail on specific products and miss the emissions from smaller, niche activities, but offers the warranty that all emissions are taken into account (Proposition 1).
Lifecycle Analysis estimates a specific product’s environmental impact over its entire lifecycle, from raw material extraction to disposal. However, the under or over counting of secondary emissions are only estimated and the total complexity makes a consistency check impossible to actually carry out.
A concrete question
@@
To estimate the carbon footprint of a T-shirt for instance, one needs the footprint of the cotton all the way from production of the raw material up to the stitching of the T-shirt, the washing during its lifespan and the disposal of the unusable T-shirt. But to estimate the cotton footprint only, one needs the transport of the water, the production of the pesticides to protect the crop, the fertilizers, all those requiring tractors and fuel, the fuel itself requiring transport and fuel, thus creating loops that are difficult to extricate. And once estimated all the emissions, how do we know that all the emissions of all the objects that we so compute, actually sum up to all the emissions that we actually produce? The Input-Output method takes the problem from the other end: it takes all the measured emissions, and then attributes it to different sectors, and then computes the carbon footprint of a goods from the textile industry according to its direct emissions, combined with all the indirect emissions as we now explain.
Basic Input-Output
@@
[…]
@@
for $i=1,\dots,n$. The value $t_{i}$ represents the total output of the sector $S_{i}$, which is also the total input because the system is a closed circuit.
Definition 1.
The technical coefficient matrix $A=(a_{ij})$ is given by
$a_{ij}=\frac{c_{ij}}{t_{j}},$
so that $A=CD_{T}^{-1}$ where $D_{T}$ denotes the diagonal matrix with entries $t_{1},\dots,t_{n}$ in the diagonal.
Recall that the scalar product of two vectors $X=(xi,\dots,x_{n})$ and $Y^{t}=(y_{1},\dots,y_{n})$ is given by
$\left<X,Y\right>=\sum_{i=1}^{n}x_{i}y_{i}=XY$
where $X$ is a line vectors, namely $1\times n$ matrix, and $Y$ is a column vector, namely an $n\times 1$ matrix and XYXY is the usual matrix multiplication so a real number.
Footprints
Economical activities have all sorts of footprints that are correlated to the income and the cost of things. To compute the footprint we need to have a emission vector $E=(e_{1},\dots,e_{n})$ where $e_{i}$ is the direct emission of sector $S_{i}$. It could be millions of tons of $CO_{2}$, measured or estimated directly above the industries of the sector, or square feet measured by the occupancy of the industries, or tons of plastic trash estimated by the collectivities, etc. Then $|E|=e_{1}+\cdots+e_{n}$ is the total footprint of the economical system. The direct intensity is given by $F=(f_{1},\dots,f_{n})$ with
@@
$\left<D,F\right>=\sum_{i=1}^{n}d_{i}f_{i}$
it is the footprint directly attributed to the consumers, but it doesn’t take into account all the intermediate exchanges that the sectors had to create the product. If you buy goods for $k$ dollars to industry $S_{i}$, then $kf_{i}$ indicates the direct footprint of the purchase, but not the total footprint of the goods. That one has to take into account the exchanges the sector $S_{i}$ had from the other ones.
Definition 2.
The total intensity vector $X=(x_{1},\cdots,x_{n})$ giving the total footprint of the sector $S_{i}$ by money amount of final consumer demand, is given by
@@
where II denotes the $n\times n$ identity matrix and AA the technical coefficients matrix from Definition 1.
Notice that since most matrices are invertible we’$l$ assume that this one is.
@@
The total footprint is completely attributed to the consumers. Namely, with the above notations, $\left<X,D\right>=|E|$.
@@
First notice that the equations $t_{i}=\sum_{j=1}^{n}c_{ij}+d_{i}$ for $i=1,\dots,n$ rephrase as
$D=T-C{\bf 1}=T-AD_{T}{\bf 1}=T-AT=(I-A)T,$
where ${\bf 1}=(1,\dots,1)$ and $T$ is the total vector and $D$ the demand vector, both viewed as column vectors. We compute:
$\left<X,D\right>=\left<F(I-A)^{-1},(I-A)T\right>=FT=\sum_{i=1}^{n}\frac{e_{i}}{t_{i}}t_{i}=|E|$
@@
Corporations
The carbon footprint inventory of a specific firm or corporation is separated in its direct emissions or scope one, and its indirect emissions. The latter are separated as scope two consisting of the emissions from acquired electricity, heat, and cooling, and the rest is scope three, computed using databases of lifecycle analysis [GHG].
Inverses
If $B$ is a matrix $\epsilon$-close to AA (say in terms of the coefficients), then the distance between $(I-A)^{-1}$ and $(I-B)^{-1}$ could a priori be as big as we want, for instance if the matrices both have an eigenvalue close enough to 1. Since all the values are estimated, the model is probably not very stable in an optimization exercise.
@@
The consumer footprint doesn’t take into account the added value $V$, which is what the companies make, and the model attributes the footprint solely to the consumers, through the vector $D$. We could define the total systemic vector
$Y=F(I-A^{t})^{-1}$
and check that $\left<Y,V\right>=|E|$. That vector distributes the total footprint to the added value of the different sectors. That attribution seems more natural than the consumer one, as it links the footprint to the benefits of the companies, see also the 2021 opinion piece [NYTop].
Concretely
The tables are quite available, see for Switzerland [IOCH], but a measure of environmental impact for each sector appearing in the table is harder to obtain, especially as it needs to be computed/collected in an homogeneous manner over the different sectors in order to make sense.
Acknowledgements
This note is based on the article by J. Kitzes [JKit], the talk by J. Steinberger at the Climathics 2022 conference [Clim], and the third author’s talk for the Comission Romande de mathématiques in 2024.